# Oryx: a Performant and Scalable Algorithm for Many-Agent Coordination in Offline MARL

### v1

@@ Abstract
A key challenge in offline multi-agent reinforcement learning (MARL) is achieving effective many-agent multi-step coordination in complex environments. In this work, we propose Oryx, a novel algorithm for offline cooperative MARL to directly address this challenge. Oryx adapts the recently proposed retention-based architecture Sable (Mahjoub et al., 2025) and combines it with a sequential form of implicit constraint Q-learning (ICQ) (Yang et al., 2021), to develop a novel offline auto-regressive policy update scheme. This allows Oryx to solve complex coordination challenges while maintaining temporal coherence over lengthy trajectories. We evaluate Oryx across a diverse set of benchmarks from prior works—SMAC, RWARE, and Multi-Agent MuJoCo—covering tasks of both discrete and continuous control, varying in scale and difficulty. Oryx achieves state-of-the-art performance on more than 80% of the 65 tested datasets, outperforming prior offline MARL methods and demonstrating robust generalisation across domains with many agents and long horizons. Finally, we introduce new datasets to push the limits of many-agent coordination in offline MARL, and demonstrate Oryx’s superior ability to scale effectively in such settings. We will make all of our datasets, experimental data, and code available upon publication.
@@ 1 Introduction
Offline MARL aims to address this exact challenge, training multi-agent policies solely from pre-collected data without further environment interaction. However, learning in the offline multi-agent setting introduces two primary difficulties. The first, well-studied challenge is accumulating extrapolation error, which occurs when agents select actions during training that fall outside the distribution of the offline dataset. This issue compounds rapidly as the joint action space grows exponentially with the number of agents, (Yang et al., 2021). Recent works (Wang et al., 2023; Matsunaga et al., 2023; Shao et al., 2023; Bui et al., 2025) have made progress on addressing extrapolation error through policy constraints or conservative value estimation. However, such methods were typically only tested on settings with relatively few agents. Leaving open the question of whether such methods are able to scale to more agents.
@@
To jointly tackle the two fundamental challenges of extrapolation error and miscoordination in many-agent settings, we propose Oryx, a novel offline MARL algorithm that unifies scalable sequence modelling with constrained offline policy improvement and structured sequential coordination. Oryx integrates the retention-based sequence modelling architecture of Sable (Mahjoub et al., 2025) with an enhanced offline multi-agent objective based on ICQ (Yang et al., 2021). Furthermore, by leveraging a dual-decoder architecture—simultaneously predicting actions and Q-values—Oryx is able to use a counterfactual advantage (Foerster et al., 2018), enabling robust and extrapolation-safe policy updates. Finally, Oryx’s sequential policy updating scheme explicitly addresses miscoordination by conditioning each agent’s policy update on the actions already executed by other agents in sequence, thus ensuring stable policy improvement.
We extensively evaluate Oryx across a broad set of challenging benchmarks—SMAC, RWARE and Multi-Agent MuJoCo—that include discrete and continuous control, varying episode lengths, and diverse agent densities. Oryx sets a new state-of-the-art in offline MARL, outperforming existing approaches in more than 80% of the 65 evaluated datasets. Furthermore, we specifically test Oryx in large many-agent settings. To achieve this, we create new datasets (with up to 50 agents) and show that Oryx maintains its superior performance at such scales. We make all of our datasets, experimental data, and code available upon publication.
Overall, by addressing both extrapolation error and miscoordination in a unified and scalable framework, Oryx significantly advances offline MARL, bringing us closer to being able to reliably deploy cooperative, multi-agent policies learned entirely from static data in complex, real-world domains.
@@ 2 Background
In our work, we adopt the multi-agent notation from Zhong et al. (2024b). Specifically, let $i_{1:m}$ denote an ordered subset $\{i_{1},\dots,i_{m}\}$ of $\mathcal{N}$, then $-i_{1:m}$ refers to its complement, and for $m=1$, we have $i$ and $-i$, respectively. The $k^{th}$ agent in the ordered subset is indexed as $i_{k}$. The action-value function is then defined as
@@
Implicit constraint Q-learning (ICQ) – effective offline regularisation. The key idea in ICQ (Yang et al., 2021) is to avoid out-of-distribution actions in the offline setting by computing target Q-values sampled from the behaviour policy $\mu$ (extracted from the dataset), instead of $\pi$, such that the Bellman operator becomes $(\mathcal{T}^{\pi}Q)(\tau,a)=r+\gamma\mathbb{E}_{a^{\prime}\sim\mu}[\rho(\tau^{\prime},a^{\prime})Q(\tau^{\prime},a^{\prime})]$, where $\rho(\tau^{\prime},a^{\prime})=\frac{\pi(a^{\prime}|\tau^{\prime})}{\mu(a^{\prime}|\tau^{\prime})}$ is an importance sampling weight. However, obtaining an accurate $\mu$ is itself difficult. Therfore, ICQ instead constrains policy updates such that $D_{KL}(\pi\parallel\mu)[\tau]\leq\epsilon$, with the corresponding optimal policy taking the form $\pi_{k+1}^{*}(a|\tau)=\frac{1}{Z(\tau)}\mu(a|\tau)\exp\left(\frac{Q_{\pi_{k}}(\tau,a)}{\alpha}\right)$. Here, $\alpha>0$ is a Lagrangian coefficient in the unconstrained optimisation objective and $Z(\tau)=\sum_{\tilde{a}}\mu(\tilde{a}|\tau)\exp\left(\frac{1}{\alpha}Q_{\pi_{k}}(\tau,\tilde{a})\right)$ is the normalisation function. Finally, solving for $\rho(\tau^{\prime},a^{\prime})=\frac{\pi_{k+1}^{*}(a^{\prime}|\tau^{\prime})}{\mu(a^{\prime}|\tau^{\prime})}$, gives the ICQ operator (that only requires sampling from $\mu$) as
@@ 3 Method
Oryx integrates the strengths of Sable’s auto-regressive retention-based architecture and ICQ’s robust regularisation to address the specific challenges of long-horizon miscoordination and accumulating extrapolation error in offline learning. We begin by deriving an ICQ regularised sequential updating scheme.
A particularly effective model class that can leverage the advantage decomposition from Kuba et al. (2021) is sequence models, with notable examples the multi-agent transformer (MAT) (Wen et al., 2022) and Sable (Mahjoub et al., 2025). These models naturally represent the joint policy as a product of factors that similarly decompose auto-regressively as $\pi(\boldsymbol{a}|\boldsymbol{\tau})=\prod^{n}_{j=1}\pi^{i_{j}}(a^{i_{j}}|\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}})$. We use this fact and the advantage decomposition theorem to derive Theorem 1 (for proof see Appendix).
@@
For an auto-regressive model, the multi-agent joint-policy under ICQ regularisation can be optimised sequentially for $j=1,...,n$ over a data batch $\mathcal{B}$ as follows:
@@
Finally, the centralised advantage estimate in the multi-agent policy gradient is susceptible to high variance. In particular, Kuba et al. (2021) provides an upper bound on the difference in gradient variance between independent and centralised learning when using the standard V-value function as a baseline as $(n-1)\frac{(\epsilon B_{i})^{2}}{1-\gamma^{2}}$, where $B_{i}=\sup_{\boldsymbol{\tau},\mathbf{a}}||\nabla_{\theta^{i}}\log\pi_{\theta}^{i}(\hat{a}^{i}|\boldsymbol{\tau})||$, $\epsilon_{i}=\sup_{\boldsymbol{\tau},\mathbf{a}^{-i},a^{i}}|A^{i}(\boldsymbol{\tau},\mathbf{a}^{-i},a^{i})|$, and $\epsilon=\max_{i}\epsilon_{i}$. This bound grows linearly in the number of agents. We can arrive at a better bound that removes the $(n-1)$ term by employing a counterfactual baseline (Kuba et al., 2021) as used in (Foerster et al., 2018) such that
@@
This completes the sequential updating scheme for Oryx. A full algorithm description is provided in Algorithm 1. Next, we describe the architectural details allowing Oryx to model long-term dependencies across sampled trajectories from the dataset.
Algorithm 1 Oryx’s sequential updating scheme with ICQ regularisation
@@
3:     Sample a mini-batch =$\mathcal{B}=\{(\boldsymbol{\tau},\mathbf{a}^{i_{1:j}},\boldsymbol{\tau}^{\prime},\mathbf{a}^{\prime,i_{1:j}})\}$ of trajectories from the offline dataset $\mathcal{D}$.
@@ 4 Results
In this section, we conduct detailed empirical evaluations to substantiate the core contributions of our proposed algorithm, Oryx. We begin by rigorously validating its key design components: (i) sequential action selection for agent coordination, (ii) a memory mechanism for temporal coherence, and (iii) autoregressive ICQ for stable offline training. To assess scalability, we subject Oryx to a demanding multi-agent setting involving up to 50 agents. Specifically, we use the Connector environment (Bonnet et al., 2024), which progressively increases coordination complexity as the number of agents grows. Finally, we perform an extensive and diverse benchmarking study across more than 65 datasets from prominent MARL benchmarks, covering a wide range of tasks in SMAC (Samvelyan et al., 2019b), MAMuJoCo (Peng et al., 2021), and RWARE (Papoudakis et al., 2021).
(a) Visualisation of ’T’-Maze
@@
(b) Results across baselines and ablations
Figure 2: Evaluating long horizon coordination.(a) T-Maze layout: A minimal two-agent environment where the target is revealed only at the first timestep, requiring agents to retain goal information throughout the episode. (b) Performance: Oryx successfully solves the task only when all components are present, while baseline methods fail to perform across both dataset types.
@@ 4.1 Validating Oryx’s core mechanisms
Inspired by Osband et al. (2020), we designed a ’T’-Maze environment (see Figure 2(a)) to specifically isolate a long-horizon multi-agent coordination challenge. On the first timestep, two agents each independently select a colour ("green" or "orange"). On the subsequent timestep, they spawn randomly at the bottom of the maze and briefly observe their choice of goal action and the goal locations (e.g., orange-left, green-right). Without further observing their goal, agents must navigate to their respective targets, manoeuvring at the junction to avoid collision and delay. Success requires: (i) selecting different colours initially for effective coordination, (ii) retaining memory of their goal choice action, and (iii) efficiently manoeuvring around one another. Agents receive a team reward of 1 if both agents reach the goal.
We generated two datasets: a mixed dataset containing primarily unsuccessful trajectories and a smaller number of successful examples, and an expert dataset consisting solely of successful trajectories. Dataset statistics are detailed in the appendix following Formanek et al. (2024a). We evaluated several baselines: i) fully independent ICQ learners (I-ICQ), ii) the CTDE variant with mixing networks (MAICQ (Yang et al., 2021)), and iii) targeted Oryx ablations disabling sequential action selection, memory, and offline autoregressive ICQ regularisation individually to isolate their contributions.
As shown in Figure 2(b), Oryx consistently achieves optimal performance with both datasets, while baseline ICQ variants fail regardless of value decomposition. The ablation results confirm that each of Oryx’s core components—sequential action selection, memory, and ICQ-based offline regularisation—is individually crucial for enabling effective long-horizon coordination among agents.
@@ 4.2 Testing coordination in complex many-agent settings
Having validated the importance of Oryx’s core components in a smaller-scale setting, we now stress-test the algorithm in significantly larger and more challenging many-agent coordination scenarios. To effectively evaluate this, we select the Connector environment from Bonnet et al. (2024), which inherently becomes more complex as agent density increases. It is important to highlight that naively increasing agent numbers does not necessarily enhance task complexity; it may even simplify certain problems where coordination is less critical. Despite its popularity, many SMAC scenarios tend to become easier as the number of agents increase (for example, reported performance often takes on the following task order, 5m_vs_6m $>$ 8m_vs_9m $>$ 10m_vs_11m (Wen et al., 2022; Yu et al., 2022; Hu et al., 2023)). In Connector, agents spawn randomly on a fixed-size grid and are each assigned a random target location. Agents must navigate to their targets, leaving impassable trails behind them that can obstruct other agents. The necessity for careful coordination sharply increases as agent density grows, making Connector ideal for testing Oryx.
Datasets for Connector were generated by recording replay data from online training sessions using Sable (Mahjoub et al., 2025). Due to the substantial volume of data generated by the online systems (20M timesteps), we applied random uniform subsampling on each dataset to select 1M timesteps. Statistical summaries of these datasets are provided in the appendix. For comparison, we again use MAICQ (Yang et al., 2021) as our baseline. MAICQ is particularly valuable as a comparative algorithm because it incorporates several widely used components that differ notably from Oryx, specifically: (i) RNN-based memory, (ii) global state-conditioned value decomposition, and (iii) the original non-autoregressive ICQ loss. We normalise the results as in Fu et al. (2020) $norm\_score=\frac{score-random\_score}{expert\_score-random\_score}$, where $random\_score$ is the average score achieved by agents taking random actions and the $expert\_score$ is the performance of the online system at the end of training.
@@
Figure 3: Evaluating Oryx on Connector with varying number of agents.
@@ 4.3 Demonstrating state-of-the-art performance on existing offline MARL datasets
Figure 4: Performance of Oryx across diverse benchmark datasets from prior literature. Scores are normalised relative to the current state-of-the-art, with values above 1 indicating that Oryx surpasses previous best-known results. Unnormalized scores are provided in the appendix. Gold stars indicate instances where Oryx matches or exceeds state-of-the-art performance, while black stars denote otherwise. Statistical significance was determined using a simple two-sided t-test (Papoudakis et al., 2021). Oryx results are averaged over 10 random seeds, each evaluated at the end of training over 320 episodes. Baseline algorithm results are directly sourced from their respective original publications, consistent with the approach of Formanek et al. (2024b).

### v2

# Oryx: a Scalable Sequence Model for Many-Agent Coordination in Offline MARL

@@ Abstract
A key challenge in offline multi-agent reinforcement learning (MARL) is achieving effective many-agent multi-step coordination in complex environments. In this work, we propose Oryx, a novel algorithm for offline cooperative MARL to directly address this challenge. Oryx adapts the recently proposed retention-based architecture Sable (Mahjoub et al., 2025) and combines it with a sequential form of implicit constraint Q-learning (ICQ) (Yang et al., 2021), to develop a novel offline autoregressive policy update scheme. This allows Oryx to solve complex coordination challenges while maintaining temporal coherence over long trajectories. We evaluate Oryx across a diverse set of benchmarks from prior works—SMAC, RWARE, and Multi-Agent MuJoCo—covering tasks of both discrete and continuous control, varying in scale and difficulty. Oryx achieves state-of-the-art performance on more than 80% of the 65 tested datasets, outperforming prior offline MARL methods and demonstrating robust generalisation across domains with many agents and long horizons. Finally, we introduce new datasets to push the limits of many-agent coordination in offline MARL, and demonstrate Oryx’s superior ability to scale effectively in such settings.
@@ 1 Introduction
Offline MARL aims to address this exact challenge, training multi-agent policies solely from pre-collected data without further environment interaction. However, learning in the offline multi-agent setting introduces two primary difficulties. The first, well-studied challenge is accumulating extrapolation error, which occurs when agents select actions during training that fall outside the distribution of the offline dataset. This issue compounds rapidly as the joint action space grows exponentially with the number of agents (Yang et al., 2021). Recent works (Wang et al., 2023; Matsunaga et al., 2023; Shao et al., 2023; Bui et al., 2025) have made progress on addressing extrapolation error through policy constraints or conservative value estimation. However, such methods were typically only tested on settings with relatively few agents. Leaving open the question of whether such methods are able to scale to more agents.
@@
To jointly tackle the two fundamental challenges of extrapolation error and miscoordination in many-agent settings, we propose Oryx, a novel offline MARL algorithm that unifies scalable sequence modelling with constrained offline policy improvement. Oryx integrates the retention-based sequence modelling architecture of Sable (Mahjoub et al., 2025) with an enhanced offline multi-agent objective based on ICQ (Yang et al., 2021). Furthermore, by leveraging a dual-decoder architecture—simultaneously predicting actions and Q-values—Oryx is able to use a counterfactual advantage (Foerster et al., 2018), enabling robust and extrapolation-safe policy updates. Finally, Oryx’s sequential policy updating scheme explicitly addresses miscoordination by conditioning each agent’s policy update on the actions already executed by other agents in sequence, thus ensuring stable policy improvement.
We extensively evaluate Oryx across a broad set of challenging benchmarks—SMAC, RWARE and Multi-Agent MuJoCo—that include discrete and continuous control, varying episode lengths, and diverse agent densities. Oryx sets a new state-of-the-art in offline MARL, outperforming existing approaches in more than 80% of the 65 evaluated datasets. Furthermore, we specifically test Oryx in large many-agent settings. To achieve this, we create new datasets (with up to 50 agents) and show that Oryx maintains its superior performance at such scales. Overall, by addressing both extrapolation error and miscoordination in a scalable sequence modelling framework, Oryx significantly advances offline MARL, bringing us closer to being able to reliably deploy cooperative, multi-agent policies learned entirely from static data in complex, real-world domains. To accelarete future research in this direction, we make all of our datasets and code available on GitHub [^1] https://github.com/instadeepai/og-marl.
@@ 2 Background
In our work, we adopt the multi-agent notation from Zhong et al. (2024b). Specifically, let $i_{1:m}$ denote an ordered subset $\{i_{1},\dots,i_{m}\}$ of $\mathcal{N}$, then $i_{1:m}$ refers to its complement, and for $m=1$, we have $i$ and i-i, respectively. The $k^{th}$ agent in the ordered subset is indexed as $i_{k}$. The action-value function is then defined as
@@
[…]
@@
on pairs present within the dataset. It does so by formulating policy updates under an implicit constraint optimization framework, ensuring the learned policy closely follows the behavior policy encoded in the data.
@@ 3 Method
Oryx integrates the strengths of Sable’s auto-regressive retention-based architecture and ICQ’s robust offline regularisation to address the specific challenges of long-horizon miscoordination and accumulating extrapolation error in offline MARL. A particularly effective model class that can naturally instantiate autoregressive policies is sequence models, with notable examples the multi-agent transformer (MAT) (Wen et al., 2022) and Sable (Mahjoub et al., 2025). These models naturally represent the joint policy as a product of factors that similarly decompose auto-regressively as $\pi(\boldsymbol{a}|\boldsymbol{\tau})=\prod^{n}_{j=1}\pi^{i_{j}}(a^{i_{j}}|\boldsymbol{\tau},\mathbf{a}^{i_{1:j-1}})$.
Autoregressive ICQ loss. Next, we use the Oryx network’s autoregressive policy structure and the advantage decomposition theorem  (Kuba et al., 2021) to derive Theorem 1 (for proof see Appendix).
@@
For an auto-regressive model, the multi-agent joint-policy under ICQ regularisation can be optimised sequentially for j=1,,nj=1,...,n over a data batch $\mathcal{B}$ as follows:
@@
Finally, the centralised advantage estimate in the multi-agent policy gradient is susceptible to high variance. In particular, Kuba et al. (2021) provides an upper bound on the difference in gradient variance between independent and centralised learning when using the standard V-value function as a baseline as $(n-1)\frac{(\epsilon B_{i})^{2}}{1-\gamma^{2}}$, where $B_{i}=\sup_{\boldsymbol{\tau},\mathbf{a}}||\nabla_{\theta^{i}}\log\pi_{\theta}^{i}(\hat{a}^{i}|\boldsymbol{\tau})||$, $\epsilon_{i}=\sup_{\boldsymbol{\tau},\mathbf{a}^{-i},a^{i}}|A^{i}(\boldsymbol{\tau},\mathbf{a}^{-i},a^{i})|$, and $\epsilon=\max_{i}\epsilon_{i}$. This bound grows linearly in the number of agents. We can arrive at a better bound that removes the (n1)(n-1) term by employing a counterfactual baseline (Kuba et al., 2021) as used in Foerster et al. (2018) such that
@@
This completes the sequential updating scheme for Oryx. A full algorithm description is provided in 1. Next, we describe the architectural details allowing Oryx to model long-term dependencies across sampled trajectories from the dataset.
Algorithm 1 Oryx’s sequential updating scheme with autoregressive ICQ regularisation
@@
3:  Sample a mini-batch $\mathcal{B}=\{(\boldsymbol{\tau},\mathbf{a}^{i_{1:j}},\boldsymbol{\tau}^{\prime},\mathbf{a}^{\prime,i_{1:j}})\}$ of trajectories from the offline dataset $\mathcal{D}$.
@@ 4 Results
In this section, we conduct detailed empirical evaluations to substantiate the core contributions of our proposed algorithm, Oryx. We begin by rigorously validating its key design components: (i) sequential action selection for agent coordination, (ii) a memory mechanism for temporal coherence, and (iii) autoregressive ICQ for stable offline training. To assess scalability, we subject Oryx to a demanding multi-agent setting involving up to 50 agents. Specifically, we use the Connector environment (Bonnet et al., 2024), which progressively increases coordination complexity as the number of agents grows. Finally, we perform an extensive and diverse benchmarking study across more than 65 datasets from prominent MARL benchmarks, covering a wide range of tasks in SMAC (Samvelyan et al., 2019), MAMuJoCo (Peng et al., 2021), and RWARE (Papoudakis et al., 2021).
(a) T-Maze
@@
(b) Performance across baselines and ablations
Figure 2: Evaluating long horizon coordination. To issolate the importance of the different components of Oryx a minimal two-agent environment, T-Maze was designed. In the environment the target states are revealed only at the first timestep, requiring agents to retain goal information throughout the episode and carefully coordinate at the end. Oryx successfully solves the task only when all components are present, while baseline methods fail to perform across both the replay and expert datasets.
@@ 4.1 Validating Oryx’s core mechanisms
Inspired by Osband et al. (2020), we designed a T-Maze environment (see Figure 2) to specifically isolate a long-horizon multi-agent coordination challenge. On the first timestep, two agents each independently select a colour ("green" or "orange"). On the subsequent timestep, they spawn randomly at the bottom of the maze and briefly observe their choice of goal action and the goal locations (e.g., orange-left, green-right). Without further observing their goal, agents must navigate to their respective targets, manoeuvring at the junction to avoid collision and delay. Success requires: (i) selecting different colours initially for effective coordination, (ii) retaining memory of their goal choice action, and (iii) efficiently manoeuvring around one another. Agents receive a team reward of 1 if both agents reach the goal. We generated two datasets: a mixed dataset containing primarily unsuccessful trajectories and a smaller number of successful examples, and an expert dataset consisting solely of successful trajectories. Dataset statistics are detailed in the appendix following Formanek et al. (2024a). We evaluated several baselines: i) fully independent ICQ learners (I-ICQ), ii) the CTDE variant with mixing networks (MAICQ (Yang et al., 2021)), and iii) targeted Oryx ablations disabling autoregressive action selection, memory, and offline sequential ICQ regularisation individually to isolate their contributions.
As shown in Figure 2, Oryx consistently achieves optimal performance with both datasets, while baseline ICQ variants fail regardless of value decomposition. The ablation results confirm that each of Oryx’s core components—sequential action selection, memory, and ICQ-based offline regularisation—is individually crucial for enabling effective long-horizon coordination among agents.
@@ 4.2 Testing coordination in complex many-agent settings
Having validated the importance of Oryx’s core components in a smaller-scale setting, we now stress-test the algorithm in significantly larger and more challenging many-agent coordination scenarios. To effectively evaluate this, we select the Connector environment from Bonnet et al. (2024), which inherently becomes more complex as agent density increases. It is important to highlight that naively increasing agent numbers does not necessarily enhance task complexity; it may even simplify certain problems where coordination is less critical. Despite its popularity, many SMAC scenarios tend to become easier as the number of agents increase (for example, reported performance often takes on the following task order, 5m_vs_6m >> 8m_vs_9m >> 10m_vs_11m (Wen et al., 2022; Yu et al., 2022; Hu et al., 2023)). In Connector, agents spawn randomly on a fixed-size grid and are each assigned a random target location. Agents must navigate to their targets, leaving impassable trails behind them that can obstruct other agents. The necessity for careful coordination sharply increases as agent density grows, making Connector ideal for testing Oryx.
Datasets for Connector were generated by recording replay data from online training sessions using Sable (Mahjoub et al., 2025). Due to the substantial volume of data generated by the online systems (20M timesteps), we applied random uniform subsampling on each dataset to select 1M timesteps. Statistical summaries of these datasets are provided in the appendix. For comparison, we again use MAICQ (Yang et al., 2021) as our baseline. MAICQ is particularly valuable as a comparative algorithm because it incorporates several widely used components that differ notably from Oryx, specifically: (i) RNN-based memory, (ii) global state-conditioned value decomposition, and (iii) the original non-autoregressive ICQ loss. We normalise the results as in Fu et al. (2020) $norm\_score=\frac{score-random\_score}{expert\_score-random\_score}$, where random_scorerandom\_score is the average score achieved by agents taking random actions and the expert_scoreexpert\_score is the performance of the online system at the end of training.
@@
Figure 3: Evaluating Oryx on many-agent settings. We compare Oryx, with its autoregressive ICQ loss and sequence model architecture, to MAICQ which is a non-autoregressive CTDE algorithm. The two algorithms are trained on datasets from Connector (Bonnet et al., 2024) scenarios with increasing numbers of agents. While the performance of MAICQ dramatically degrades on scenarios with large numbers of agents, Oryx’s performance remains robust.
@@ 4.3 Demonstrating state-of-the-art performance on existing offline MARL datasets
Figure 4: Performance of Oryx across diverse benchmark datasets from prior literature. Scores are normalised relative to the current state-of-the-art, with values above 1 indicating that Oryx surpasses previous best-known results. Unnormalized scores are provided in the appendix. Gold stars indicate instances where Oryx matches or exceeds state-of-the-art performance, while black stars denote otherwise.
@@
4.4 Comparing auto-regressive sequence modeling architectures
Oryx extends auregressive MARL sequence modeling into the offline MARL setting by leveraging Sable (Mahjoub et al., 2025) as its network architecture backbone. While prior sequence models such as MAT (Wen et al., 2022) have demonstrated strong performance in online settings, the decision to adopt Sable as the backbone for Oryx was motivated by its superior scalability, efficiency, and stability when trained across large populations of agents.
However, to further quantify the advantages of our design choices, we conducted an ablation comparing Oryx with an offline variant of MAT using identical training procedures and the same autoregressive ICQ loss. This experiment isolates the impact of choosing Sable over MAT for Oryx’s sequence model component. Following the evaluation methodology proposed by Gorsane et al. (2022), we report aggregated metrics across all SMAC and RWARE datasets. To facilitate aggregation across scenarios with different expected episode returns, results were normalised by the highest episode return in each dataset. The results in Table 1 demonstrate that while MAT+ICQ is a strong baseline, Oryx still ourperforms it across all metrics and environments, validating our design decision to build on the Sable network architecture instead using MAT.
Table 1: Comparing Oryx and MAT+ICQ. In order to isolate the effect of using the Sable network architecture as the backbone of the Oryx network, we compare it to an offline varient of MAT that uses the same autoregressive ICQ loss as Oryx (MAT+ICQ). We report several aggregated metrics with 95% stratified bootstrap confidence intervals (Agarwal et al., 2021) across all SMAC and RWARE datasets. To Facilitate aggregation across scenarios with different episode returns, scores were first normalised by the highest return in each dataset respectivly. We see from the results that on both environments the mean, median, interquartile mean (IQM), and optimality gap of Oryx is superior to MAT+ICQ.
SMAC RWARE MAT+ICQ Oryx MAT+ICQ Oryx Median $\uparrow$ 0.71 [0.62, 0.74] 0.91 [0.88, 0.92] 0.85 [0.84, 0.86] 0.89 [0.88, 0.90] IQM $\uparrow$ 0.67 [0.66, 0.69] 0.87 [0.86, 0.88] 0.85 [0.84, 0.86] 0.89 [0.89, 0.90] Mean $\uparrow$ 0.63 [0.62, 0.64] 0.77 [0.76, 0.79] 0.84 [0.83, 0.84] 0.89 [0.88, 0.90] Optimality Gap $\downarrow$ 0.38 [0.36, 0.39] 0.23 [0.22, 0.25] 0.16 [0.16, 0.17] 0.11 [0.10, 0.12]